### v1

@@ 3.1. Shelling Order
    Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. First we show that if $i_{1}<j_{1}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>\alpha_{j}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq j_{1}$ and $i_{2}\leq\alpha_{j}$.
    Suppose $i_{1}<j_{1}$. Since $i_{1}<j_{1}=\alpha_{j}-4<\alpha_{j}\leq n-3$, we get $\alpha_{j}\nsim i_{1}$ and $\alpha_{j}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $i_{1}<j_{1}=\alpha_{j}-4$, $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Let $i_{2}>\alpha_{j}$. Since $\alpha_{j}\geq 6$, we have $2\leq j_{1}=\alpha_{j}-4<\alpha_{j}<i_{2}$. So, $j_{1}\nsim\alpha_{j}$ and $j_{1}\nsim i_{2}$. Thus $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-4$ and $i_{2}\geq\alpha_{j}+1$, it follows that $F^{\prime\prime}\notin\mathcal{D}$. Hence $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1).
    Now, let $i_{1}\geq j_{1}$ and $i_{2}\leq\alpha_{j}$. Suppose $F_{i}\in\mathcal{T}_{j_{0}}$. Then $\alpha_{j}=\alpha_{i}\neq i_{2}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{1}\geq j_{1}=j_{2}-1$ implies that $i_{1}>j_{2}$. So, we have $\alpha_{j}-3=j_{2}<i_{1}<i_{2}<\alpha_{j}$. It follows that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m+1$, we get $\alpha_{i}\in\{m-1,m\}$. We have $m-3=j_{1}\leq i_{1}<i_{2}\leq\alpha_{j}=m+1$. Therefore, $W_{n}[F_{i}^{c}]$ is disconnected implies that $\alpha_{i}=m$, $i_{1}=j_{1}=m-3$ and $i_{2}=\alpha_{j}=m+1$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
@@
    Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{2},i_{2}\}$. First we show that if $i_{1}<\alpha_{j}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>j_{2}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq\alpha_{j}$ and $i_{2}\leq j_{2}$.
    Suppose $i_{1}<\alpha_{j}$. We first show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{j}\leq n-6$, we get $j_{2}=\alpha_{j}+4\leq n-2$. We have $i_{1}<\alpha_{j}<\alpha_{j}+4=j_{2}\leq n-2$. So $j_{2}\nsim\alpha_{j}$. Now, if $i_{1}\nsim\alpha_{j}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. So, let $i_{1}\sim\alpha_{j}$. Then $i_{1}\in\{\alpha_{j}-2,\alpha_{j}-1\}$ and therefore $i_{1}\geq 2$. Thus $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $i_{1}\leq\alpha_{j}-1$ and $j_{2}=\alpha_{j}+4$, it follows that $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
    Now, suppose $i_{2}>j_{2}$. We have $4\leq\alpha_{j}<\alpha_{j}+4=j_{2}<i_{2}$. Thus $\alpha_{j}\nsim j_{2}$ and $\alpha_{j}\nsim i_{2}$. Hence, $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $i_{2}>j_{2}=\alpha_{j}+4$, $F^{\prime\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
    Finally, let $i_{1}\geq\alpha_{j}$ and $i_{2}\leq j_{2}$. Suppose $F_{i}\in\mathcal{T}_{j_{0}}$. Then $\alpha_{j}=\alpha_{i}\neq i_{1}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{2}\leq j_{2}=j_{1}+1$ implies that $i_{2}<j_{1}$. Therefore, we have $\alpha_{j}<i_{1}<i_{2}<j_{1}=\alpha_{j}+3$. This implies that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
@@
      Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $\alpha_{i}=m$. We have $m-1=\alpha_{j}\leq i_{1}<i_{2}\leq j_{2}=\alpha_{j}+4=m+3$. Therefore, $W_{n}[F_{i}^{c}]$ is disconnected implies that $i_{1}=\alpha_{j}=m-1$ and $i_{2}=j_{2}=m+3$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
@@
      Let $F^{\prime\prime\prime c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. We have $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}>m$. So by Remark 3.1 (iv), $\alpha_{j}>\alpha_{i}\geq 2m-\alpha_{j}\geq 2m-(n-6)\geq 6$. Then, we have $6\leq\alpha_{i}<\alpha_{j}<\alpha_{j}+3=j_{1}<j_{2}$. It follows that $\alpha_{i}\nsim j_{1}$ and $\alpha_{i}\nsim j_{2}$. Hence $F^{\prime\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, Proposition 3.5 implies that $F^{\prime\prime\prime}\prec F_{j}$. Hence $F_{r}=F^{\prime\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
From Lemmas 3.7, 3.8 and 3.9, it follows that for any $F_{i},F_{j}\in M(\Delta_{3}(W_{n}))$ with $F_{i}\prec F_{j}$, there exists an $F_{r}\prec F_{j}$ such that $F_{i}\cap F_{j}\subseteq F_{r}\cap F_{j}$ and $|F_{r}\cap F_{j}|=$ $|F_{j}|-1$. This means that $\prec$ provides a shelling order for $\Delta_{3}(W_{n})$.
@@ 3.2. Spanning Facets
Observe that by definition, any facet $F\in M(\Delta_{3}(W_{n}))$ is a spanning facet if and only if for each $\mu\in F$, there exists $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$.
Define a subset $\mathcal{S}$ of $M(\Delta_{3}(W_{n}))$ as follows: A facet $F\in\mathcal{S}$ if and only if there exists $s$ such that $F\in\mathcal{T}_{s}$ and $F$ satisfies one of the following conditions:
@@ 1. $(\mathcal{S}_{1})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}=3$ and $s_{1}\in V(W_{n})\setminus(\{0,1,2,3,\ldots,2m-4\}\sqcup\{n-1\})$.
@@ 2. $(\mathcal{S}_{2})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{4,5,\ldots,m-2\}$ and $s_{1}\in V(W_{n})\setminus(\{\alpha_{s}-4,\alpha_{s}-3,\alpha_{s}-2\}\sqcup\{\alpha_{s},\alpha_{s}+1,\ldots,2m-\alpha_{s}-1\}\sqcup\{n-1\})$.
@@ 3. $(\mathcal{S}_{3})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{m-1,m,m+1,m+2,\ldots,n-3\}$, $s_{1}\in V(W_{n})\setminus(\{\omega,\omega+1,\ldots,\alpha_{s},\alpha_{s}+1,\alpha_{s}+2,\alpha_{s}+3\text{ (mod $n$})\}\sqcup\{y\})$, $\omega=\min\{2m-\alpha_{s},\alpha_{s}-4\}$ and
@@
Our aim is to show that a facet $F\in M(\Delta_{3}(W_{n}))$ is a spanning facet if and only if $F\in\mathcal{S}$.
@@
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $F\in\mathcal{T}_{s}$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$ and ${F^{\prime}}^{c}=\{\alpha_{s},s_{1},\mu\}$ for some $\mu\in F$. If $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\ll F$.
@@
Since $\alpha_{s}<_{\mathcal{O}}\mu,s_{1}$, we have $F^{\prime}\in\mathcal{T}_{s}$. If $\mu<s_{1}$, then $F^{\prime c}=\{\alpha_{s}\}\sqcup\{\mu,s_{1}\}$ and $F^{\prime}\ll F$. On the other hand, if $\mu>s_{1}$, then $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s_{1},\mu\}$, and thus $\mu<n-1$ implies that $F^{\prime}\ll F$. ∎
@@
Let $F\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. Let $\mathcal{U}=V(W_{n})\setminus\bigl{(}\{\alpha_{s}-4\ (\text{mod $n$}),\alpha_{s}-3,\alpha_{s}-2,\ldots,2m-\alpha_{s}-1\}\cup\{n-1\}\bigr{)}$. Suppose for each $\alpha_{s}\in\{3,4,\ldots,m-2\}$, $s_{1}\in\mathcal{U}$. Then $F$ is a spanning facet.

### v2

@@ 3.1. Shelling Order
    Let $F^{\prime c}:=(F_{j}^{c}\setminus\{j_{2}\})\sqcup\{i_{1}\}$ and $F^{\prime\prime c}:=(F_{j}^{c}\setminus\{j_{2}\})\sqcup\{i_{2}\}$. First we show that if $i_{1}<j_{1}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>\alpha_{j}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq j_{1}$ and $i_{2}\leq\alpha_{j}$.
    Suppose $i_{1}<j_{1}$. Since $\alpha_{j}<_{\mathcal{O}}j_{1}$, $j_{1}<\alpha_{j}$ and $\alpha_{j}>m$, we get $j_{1}<2m-\alpha_{j}$ by Remark 3.1 (ii), and hence $i_{1}<2m-\alpha_{j}$, thereby implying that $\alpha_{j}<_{\mathcal{O}}i_{1}$. This means that $F^{\prime c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$. Since $i_{1}<j_{1}=\alpha_{j}-4<\alpha_{j}\leq n-3$, we get $\alpha_{j}\nsim i_{1}$ and $\alpha_{j}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Further, since $i_{1}<j_{1}=\alpha_{j}-4$, $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.5 (iii). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
    Let $i_{2}>\alpha_{j}$. Since $\alpha_{j}>m$, we get $\alpha_{j}<_{\mathcal{O}}i_{2}$. Thus, $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}<\alpha_{j}<i_{2}$ implies that $F^{\prime\prime c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Since $\alpha_{j}\geq 6$, we have $2\leq j_{1}=\alpha_{j}-4<\alpha_{j}<i_{2}$. So, $j_{1}\nsim\alpha_{j}$ and $j_{1}\nsim i_{2}$. Thus $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $j_{1}=\alpha_{j}-4$ and $i_{2}\geq\alpha_{j}+1$, it follows that $F^{\prime\prime}\notin\mathcal{D}$. Hence $F^{\prime\prime}\prec F_{j}$ by Definition 3.5 (iii). Therefore $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1).
    Now, let $i_{1}\geq j_{1}$ and $i_{2}\leq\alpha_{j}$. Suppose $F_{i}\in\mathcal{T}_{j_{0}}$. Then $\alpha_{j}=\alpha_{i}\neq i_{2}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{1}\geq j_{1}=j_{2}-1$ implies that $i_{1}>j_{2}$. So, we have $\alpha_{j}-3=j_{2}<i_{1}<i_{2}<\alpha_{j}$. It follows that $C_{n}^{2}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m+1$, we get $\alpha_{i}\in\{m-1,m\}$. We have $m-3=j_{1}\leq i_{1}<i_{2}\leq\alpha_{j}=m+1$. Therefore, $C_{n}^{2}[F_{i}^{c}]$ is disconnected implies that $\alpha_{i}=m$, $i_{1}=j_{1}=m-3$ and $i_{2}=\alpha_{j}=m+1$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
@@
    Let $F^{\prime c}:=(F_{j}^{c}\setminus\{j_{1}\})\sqcup\{i_{1}\}$ and $F^{\prime\prime c}:=(F_{j}^{c}\setminus\{j_{1}\})\sqcup\{i_{2}\}$. First we show that if $i_{1}<\alpha_{j}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>j_{2}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq\alpha_{j}$ and $i_{2}\leq j_{2}$.
    Suppose $i_{1}<\alpha_{j}$. We first show that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $\alpha_{j}\leq n-6$, we get $j_{2}=\alpha_{j}+4\leq n-2$. We have $i_{1}<\alpha_{j}<\alpha_{j}+4=j_{2}\leq n-2$. So $j_{2}\nsim\alpha_{j}$. Now, if $i_{1}\nsim\alpha_{j}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. So, let $i_{1}\sim\alpha_{j}$. Then $i_{1}\in\{\alpha_{j}-2,\alpha_{j}-1\}$ and therefore $i_{1}\geq 2$. Thus $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. We now show that $F^{\prime}\notin\mathcal{D}$. If $\alpha_{j}<_{\mathcal{O}}i_{1}$, then $i_{1}<\alpha_{j}<j_{2}$ and $\alpha_{j}<_{\mathcal{O}}j_{2}$ implies that $F^{\prime c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Therefore, since $i_{1}\leq\alpha_{j}-1$ and $j_{2}=\alpha_{j}+4$, it follows that $F^{\prime}\notin\mathcal{D}$. Moreover, if $i_{1}<_{\mathcal{O}}\alpha_{j}$, then $\alpha_{j}<_{\mathcal{O}}j_{2}$ implies that $i_{1}<_{\mathcal{O}}j_{2}$. Thus $F^{\prime c}=\{i_{1}\}\sqcup\{\alpha_{j},j_{2}\}$. Since $j_{2}=\alpha_{j}+4>i_{1}+4$, we get $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.5 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
    Now, suppose $i_{2}>j_{2}$. If $\alpha_{j}<m$, then $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}>\alpha_{j}$ implies that $j_{2}\geq 2m-\alpha_{j}$ by Remark 3.1 (i). Hence $i_{2}>2m-\alpha_{j}$, thereby implying that $\alpha_{j}<_{\mathcal{O}}i_{2}$. Also, if $\alpha_{j}\geq m$, then $i_{2}>j_{2}>\alpha_{j}$ implies that $\alpha_{j}<_{\mathcal{O}}i_{2}$. This means that $F^{\prime c}=\{\alpha_{j}\}\sqcup\{j_{2},i_{2}\}$. We have $4\leq\alpha_{j}<\alpha_{j}+4=j_{2}<i_{2}$. Thus $\alpha_{j}\nsim j_{2}$ and $\alpha_{j}\nsim i_{2}$. Hence, $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $i_{2}>j_{2}=\alpha_{j}+4$, $F^{\prime\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.5 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
    Finally, let $i_{1}\geq\alpha_{j}$ and $i_{2}\leq j_{2}$. Suppose $F_{i}\in\mathcal{T}_{j_{0}}$. Then $\alpha_{j}=\alpha_{i}\neq i_{1}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{2}\leq j_{2}=j_{1}+1$ implies that $i_{2}<j_{1}$. Therefore, we have $\alpha_{j}<i_{1}<i_{2}<j_{1}=\alpha_{j}+3$. This implies that $C_{n}^{2}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
@@
      Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $\alpha_{i}=m$. We have $m-1=\alpha_{j}\leq i_{1}<i_{2}\leq j_{2}=\alpha_{j}+4=m+3$. Therefore, $C_{n}^{2}[F_{i}^{c}]$ is disconnected implies that $i_{1}=\alpha_{j}=m-1$ and $i_{2}=j_{2}=m+3$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
@@
      Let $F^{\prime\prime\prime c}:=(F_{j}^{c}\setminus\{\alpha_{j}\})\sqcup\{\alpha_{i}\}$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, we get $\alpha_{i}<_{\mathcal{O}}j_{1},j_{2}$. Thus $F^{\prime\prime\prime c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. We have $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}>m$. So by Remark 3.1 (iv), $\alpha_{j}>\alpha_{i}\geq 2m-\alpha_{j}\geq 2m-(n-6)\geq 6$. Then, we have $6\leq\alpha_{i}<\alpha_{j}<\alpha_{j}+3=j_{1}<j_{2}$. It follows that $\alpha_{i}\nsim j_{1}$ and $\alpha_{i}\nsim j_{2}$. Hence $F^{\prime\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, Proposition 3.6 implies that $F^{\prime\prime\prime}\prec F_{j}$. Hence $F_{r}=F^{\prime\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
From Lemmas 3.8, 3.9 and 3.10, it follows that for any $F_{i},F_{j}\in M(\Delta_{3}(C_{n}^{2}))$ with $F_{i}\prec F_{j}$, there exists an $F_{r}\prec F_{j}$ such that $F_{i}\cap F_{j}\subseteq F_{r}\cap F_{j}$ and $|F_{r}\cap F_{j}|=$ $|F_{j}|-1$. This means that $\prec$ provides a shelling order for $\Delta_{3}(C_{n}^{2})$.
@@ 3.2. Spanning Facets
Observe that by definition, any facet $F\in M(\Delta_{3}(C_{n}^{2}))$ is a spanning facet if and only if for each $\mu\in F$, there exists $\tilde{F}\in M(\Delta_{3}(C_{n}^{2}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$.
Define a subset $\mathcal{S}$ of $M(\Delta_{3}(C_{n}^{2}))$ as follows: A facet $F\in\mathcal{S}$ if and only if there exists $s$ such that $F\in\mathcal{T}_{s}$ and $F$ satisfies one of the following conditions:
@@ 1. $(\mathcal{S}_{1})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}=3$ and $s_{1}\in V(C_{n}^{2})\setminus(\{0,1,2,3,\ldots,2m-4\}\sqcup\{n-1\})$.
@@ 2. $(\mathcal{S}_{2})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{4,5,\ldots,m-2\}$ and $s_{1}\in V(C_{n}^{2})\setminus(\{\alpha_{s}-4,\alpha_{s}-3,\alpha_{s}-2\}\sqcup\{\alpha_{s},\alpha_{s}+1,\ldots,2m-\alpha_{s}-1\}\sqcup\{n-1\})$.
@@ 3. $(\mathcal{S}_{3})$
    $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{m-1,m,m+1,m+2,\ldots,n-3\}$, $s_{1}\in V(C_{n}^{2})\setminus(\{\omega,\omega+1,\ldots,\alpha_{s},\alpha_{s}+1,\alpha_{s}+2,\alpha_{s}+3\text{ (mod $n$})\}\sqcup\{y\})$, $\omega=\min\{2m-\alpha_{s},\alpha_{s}-4\}$ and
@@
Our aim is to show that a facet $F\in M(\Delta_{3}(C_{n}^{2}))$ is a spanning facet if and only if $F\in\mathcal{S}$.
@@
Let $F,F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Let $F\in\mathcal{T}_{s}$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$ and ${F^{\prime}}^{c}=\{\alpha_{s},s_{1},\mu\}$ for some $\mu\in F$. If $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\in\mathcal{T}_{s}$ and $F^{\prime}\ll F$.
@@
Since $\alpha_{s}<_{\mathcal{O}}\mu,s_{1}$, we have $F^{\prime}\in\mathcal{T}_{s}$. If $\mu<s_{1}$, then $F^{\prime c}=\{\alpha_{s}\}\sqcup\{\mu,s_{1}\}$ and $F^{\prime}\ll F$ by Definition 3.4 (i). On the other hand, if $\mu>s_{1}$, then $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s_{1},\mu\}$, and thus $\mu<n-1$ implies that $F^{\prime}\ll F$ Definition 3.4 (i). ∎
@@
Let $F\in M(\Delta_{3}(C_{n}^{2}))$ such that $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. Suppose $\alpha_{s}\in\{3,4,\ldots,m-2\}$ and $s_{1}\in\mathcal{U}=V(C_{n}^{2})\setminus\bigl{(}\{\alpha_{s}-4\ (\text{mod $n$}),\alpha_{s}-3,\alpha_{s}-2,\ldots,2m-\alpha_{s}-1\}\cup\{n-1\}\bigr{)}$. Then $F$ is a spanning facet.